Find two consecutive odd positive integers such that the twice the square of the larger minus the square of the smaller is 41. Set up an algebraic equation and solve for the two numbers.
Answer by oscargut(686)

x= first number
y= second number (y=x+2)
so y is larger than x
then
2y^2-x^2=41
then
2(x+2)^2-x^2=41
then
x^2+8x-33=0
then x=[-8+sqrt(196)]/2 or x=[-8-sqrt(196)]/2
then x=3 or x=-11
then x=3 (because x and y are positives)
then
R: x=3 and y=5